OpenStudy (anonymous):

Can anyone help me with this? Let z = 1+i√3. Find the polar form of z^5 I know it's something like 2(cos ( ) isin( )) but I'm not sure what it is. is it 2(cosπ + (1/2)isinπ) = z?

OpenStudy (aravindg):

first write z in polar form apply de movries theorem for z^5 u will get the answer be happy :)

OpenStudy (anonymous):

z = 1+i√3 -> x = 1 , y = √3 => r = √ ( 1^2 + √3^2 ) = 2 , sin theta = 1/2 => theta = pi/6 Thus polar form: z = 2 cis pi/6 z^5 = 32 cis 5 * pi/6